OpenStudy (anonymous):

Designing a 11V,550mA DC power Supply: How Can I design a power supply circuit with: Input voltage: 100-240Vac Output voltage: 11Vdc , 550mA Using Full-wave Rectifier Filter and (Zener Diode circuit) Regulator in the design. Dear Friends, What should be the Ratings of the Components ?

OpenStudy (anonymous):

hi tripleA, since you have mentioned a Zener Diode Ckt regulator. what is your Vz(nearest value) what is your Vs(100-240) should be greater than Vz Imax = 550+(10% of desired o/p current) Pz > Vz * Imax R= (Vs- Vz)/ Imax Resistor power rating P > (Vs- Vz) * Imax take it from here...